RC meeting and Run Plan, Sunday November 6th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

No Meeting today

Useful Links

Current Status

Hall Status

  • Hall to CANS for laser alignment and optics tests.
    • Alignments completed during swing
    • Opportunistic coil SBS/BB OFF measurements taken. See log.
  • See Jessie's update on the configuration change here.

Open Issues

  • A final RF module needs replaced in BPM VME crate. IPM1H04A still does not send the signals we need for verification of beam position in the hall. Simona is working with Tony Delacruz on this issue and it will be addressed on Monday.
    • Simona took measurements of the signals from both IPM1H04A and IPM1H04E and compared them with the scope. See log.

Polarized Target

  • Current Polarization: N/A (Target not installed at present)
  • Hourly screenshots suspended until target installation. Lasers locked out.

Future Plans

  • Hall in restricted access for remainder of configuration change
    • We will be moving to CANS (restricted access with laser training) on Sunday swing. Check with RC if you need access during that time.
    • Please check with RC for access beyond DAQ bunker during the day

Shifts and Meetings

  • Shift leader and target operator are on stand-by today for all shifts.

Other Topics

Experimental progress (for MCC planning meeting)

  • 6.1% of accumulated charge for production data (13.55 Coulombs out of 232)
  • 30% Experimental run time (42 out of 141 days)
  • Long term "to do" items; not urgent, some can be done when 3He target not available
    • Luminosity scans with GEMs
    • Opportunistic BCM calibration (higher current) before pass change?
    • INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run
    • Misc. GEM commissioning items: plan

Run Plan After Configuration Change

Continue to update 'standard' run plan for 3-pass data taking as necessary.

Current Plan:

  • NMR measurement
  • Restore beam
  • Optics and H2 Ref Cell (Sometime during run)
  • Production Run Plan
    • 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour.
    • 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization.
    • 3.) Return to step 1.)

While on shift:

  • Keep an eye out for drifting BPM positions
  • Log all DAQ crashes in the excel worksheet along with the reason for the crash
  • If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
  • Cross check the rate with what is expected: log
  • Cross check with the expected ion chamber readings: log

Notices in Effect for Shift Crew

All shifts (TO and SL) are on standby today

Be sure to follow these procedures:

Shifts and Long Term Issues

  • Shift Openings: We're looking good for now! Schedule

Planned Accesses to the Hall

In Restricted Access. Configuration change ongoing.

Special Requests

Nothing scheduled at present